On January 1,2021,Catty Company issued$840,000 of 8%,10-year bonds for 97. Catty retired all of these bonds on January 1,2022,at 102.If Patty uses the straight-line amortization,how
much loss should be recognized on this bond retirement?
Answer #1

Par value of bonds = $840,000

Issue price =97

Cash received from issuance of bonds = Par value of bonds x Issue price

= 840,000 x 97%

= $814,800

Discount on issue of bonds = Par value of bonds- Cash received from issuance of bonds

= 840,000-814,800

= $25,200

Annual amortization of bond discount = Discount on issue of bonds / Bond life

= 25,200/10

= $2,520

Unamortized bond discount on January 1, 2022 = Discount on issue of bonds - Annual amortization of bond discount

= 25,200-2,520

= $22,680

Carrying value of bonds on January 1, 2022 = Par value of bonds - Unamortized bond discount on January 1, 2022

= 840,000-22,680

= $817,320

Cash paid to retire bonds = Par value of bonds x Call price

= 840,000 x 102%

= $856,800

Loss on bond retirement = Cash paid to retire bonds- Carrying value of bonds on January 1, 2022

= 856,800-817,320

= $39,480

Kindly comment if you need further assistance. Thanks‼!